OpenStudy (anonymous):

Write an equation for the line that passes through the points (0, -2) and (-4, 0).

OpenStudy (btaylor):

the slope between the two points is:\[\frac{rise}{run} = \frac{\Delta x}{\Delta y} = \frac{y_2-y_1}{x_2-x_1} = \frac{2}{-4} = - \frac{1}{2}\]So m = -1/2. we are given the y-intercept is b=-2. So: \[y =-\frac{1}{2} x -2\]